# OPPO: Accelerating PPO-based RLHF via Pipeline Overlap

### v1

@@ 5 Related Work
Hydra-PPO (Santacroce et al., 2023) reduces memory and latency by combining LoRA with parameter sharing across actor, critic, and reward models. Offline PPO methods (Hu et al., 2023; Noukhovitch et al., 2025) improve stability and efficiency by training from fixed preference datasets, avoiding costly online rollouts. Data-centric approaches such as LIMO (Ye et al., 2025) and S1 (Muennighoff et al., 2025) demonstrate that small, curated datasets can yield competitive performance. LIMR (Li et al., 2025) prioritizes samples using impact-based scoring, while ADARFT (Shi et al., 2025) adopts a lightweight curriculum that adjusts difficulty through reward signals. These methods primarily optimize data or optimization strategy, whereas our work focuses on improving system-level efficiency by restructuring PPO’s execution pipeline.
@@
System-level techniques seek to accelerate RLHF training by rethinking the execution stack. TRL (von Werra et al., 2020a) provides scalable multi-node training with parameter-efficient fine-tuning. OpenRLHF (Hu et al., 2025b) integrates vLLM (Kwon et al., 2023) with Ray (Liaw et al., 2018) to accelerate generation and scheduling. HybridFlow (Sheng et al., 2025) improves throughput by combining single- and multi-controller paradigms, while RLHFuse (Zhong et al., 2025) boosts GPU utilization through stage fusion and micro-batch scheduling. Our approach complements these efforts by compounding PPO’s disaggregated stages with intra- and inter-step overlap, further improving utilization and throughput.
@@ A.1 Model Evaluation Protocol
We conduct comprehensive evaluations to assess the impact of OPPO on model quality across different model scales. Our evaluation protocol compares models trained by TRL with standard PPO-based RLHF against models trained with our proposed overlapping optimization techniques. We evaluate two model configurations: Qwen2.5-7B and Qwen2.5-3B each fine-tuned on the stack exchange paired dataset for comparable training steps.
@@ A.1.4 Statistical Considerations
Evaluations are conducted on NVIDIA A100 40GB GPUs, with complete evaluation of a single model requiring approximately 2-3 hours depending on model size. The evaluation pipeline is designed to be portable and reproducible, with automated dependency management and environment configuration scripts to facilitate replication across different computational environments.

### v2

@@ 5 Related Work
Hydra-PPO (Santacroce et al., 2023) reduces memory and latency by combining LoRA with parameter sharing across actor, critic, and reward models. Offline PPO methods (Hu et al., 2023; Noukhovitch et al., 2025) improve stability and efficiency by training from fixed preference datasets, thereby avoiding costly online rollouts. Data-centric approaches such as LIMO (Ye et al., 2025) and S1 (Muennighoff et al., 2025) demonstrate that small, curated datasets can yield competitive performance. LIMR (Li et al., 2025) prioritizes samples using impact-based scoring, while ADARFT (Shi et al., 2025) adopts a lightweight curriculum that adjusts difficulty through reward signals. These methods primarily optimize data or optimization strategy, whereas our work focuses on improving system-level efficiency by restructuring PPO’s execution pipeline.
@@
System-level techniques seek to accelerate RLHF training by rethinking the execution stack. TRL (von Werra et al., 2020) provides scalable multi-node training with parameter-efficient fine-tuning. OpenRLHF (Hu et al., 2025b) integrates vLLM (Kwon et al., 2023) with Ray (Liaw et al., 2018) to accelerate generation and scheduling. HybridFlow (Sheng et al., 2025) improves throughput by combining single- and multi-controller paradigms, while RLHFuse (Zhong et al., 2025) boosts GPU utilization through stage fusion and micro-batch scheduling. Our approach complements these efforts by compounding PPO’s disaggregated stages with intra- and inter-step overlap, further improving utilization and throughput.
@@ A.1 Model Evaluation Protocol
We conduct comprehensive evaluations to assess the impact of OPPO on model quality across different model scales. Our evaluation protocol compares models trained by TRL with standard PPO-based RLHF against models trained with our proposed overlapping optimization techniques. We evaluate two model configurations: Qwen2.5-7B and Qwen2.5-3B each fine-tuned on the Stack-Exchange-Paired dataset for comparable training steps.
@@ A.1.4 Statistical Considerations
Evaluations are conducted on NVIDIA A100 40GB GPUs, with a complete evaluation of a single model requiring approximately 2-3 hours depending on model size. The evaluation pipeline is designed to be portable and reproducible, with automated dependency management and environment configuration scripts to facilitate replication across different computational environments.
A.2 The Use of Large Language Models (LLMs)
We used a large language model (ChatGPT) only as a writing assistant tool to check grammar, improve readability, and polish sentence clarity.
A.3 Reproducibility statement
To ensure reproducibility of our results, we provide comprehensive implementation details throughout the paper and supplementary materials. The complete OPPO algorithm is specified in Algorithm 1, with detailed descriptions of the intra-step and inter-step overlap mechanisms in Sections 3.1 and 3.2. Experimental configurations, including model architectures and training settings for both 3B and 7B models, are detailed in Section 4.1 and Appendix A. We utilize publicly available datasets with standard preprocessing procedures described in Section 4. The dynamic control parameters for overcommitment degree adaptation are fully specified in Sections 3.1 and 3.2, including bounds, momentum values, and window sizes. Our implementation requires minimal modifications to existing PPO codebases, with the specific integration points outlined in Section 3. All experiments were conducted on NVIDIA A100 and H200 GPUs. We will release our implementation code upon publication to facilitate reproduction and adoption of OPPO in existing RLHF pipelines.